1.     In 2012, the box industry was perfectly competitive. The lowest point on the long-run average cost curve of each of the identical box producers was $4, and this minimum point occurred at an output of 1,000 boxes per month. The market demand curve for boxes was                                                                                                     2 marks




where P was the price of a box (in dollars per box) and QD was the quantity of boxes demanded per month. The market supply curve for boxes was




Where QS was the quantity of boxes supplied per month.

a.      What was the equilibrium price of a box? Is this the long-run equilibrium price?

b.     How many firms are in this industry when it is in long-run equilibrium?




1
Expert's answer
2021-01-13T11:57:12-0500

At equilibrium:

"140000-10000P=80000+5000P \\

15000P=140000-80000 \\

15000P=60000 \\

P=\frac{60000}{15000}=4 \\

Q=140000-10000P \\

=140000-10000 \times 4 \\

=140000-40000 \\

=100000"

Minimum output per month = 1000

So no of firm "=\frac{100000}{1000}=100"

Answers:

a. Equilibrium Price = 4

b. Number of firms = 100